# PCoreSet: Effective Active Learning through Knowledge Distillation from Vision-Language Models

### v1

@@ 3.2 Our Method: ActiveKD with PCoreSet
Previous works deployed semi-supervised learning [SSL; 5, 10, 101] to leverage the large unlabeled dataset $\mathcal{D}^{(u)}=\{x_{m}^{(u)}\}_{m=1}^{M}$ at each AL round. In this paper, we deploy knowledge distillation (KD) instead of SSL thanks to the recent advances of VLMs. Specifically, we train the model for each round as following [34]:
@@
where the final loss is $\lambda\mathcal{L}_{\text{CE}}+(1-\lambda)\mathcal{L}_{\text{KD}}$, $\ell$ denotes $D_{\mathrm{KL}}$ represent cross-entropy and Kullback-Leibler divergence, respectively. For each round $r$, we use a mini-batch version of the above objective with stochastic gradient descent to optimize the parameters of $f_{r}$ by leveraging the teacher prediction $f(\cdot)$. See underlines in Alg. 1 for additional parts of ActiveKD.
@@
Figure 4: Visualization of prediction bias and its propagation. (a) Teacher hard prediction bias, (b) teacher and student soft prediction after distillation. (c) Bias propagation measured by KNN loss ($\ell_{2}$) across different numbers of clusters (k).
Foundation models such as CLIP ResNet-50 [68] inherently exhibit class imbalance [6] due to their general-purpose training objectives, as shown in Fig. 4(a). Beyond simple imbalance, this bias manifests structurally in the probability space, with predictions forming distinct clusters that occupy limited regions of the probability simplex, as visualized in Fig. 4(b). The violin plots show probability distributions of samples for given ground truth classes, revealing how teacher predictions cluster in specific regions rather than uniformly covering the probability space. Interestingly, we observe that student models inherit similar biased prediction patterns. To formalize this, we first define the teacher’s structured prediction bias as follows:
@@
In other words, all teacher predictions lie within a finite union of $\ell_{2}$-balls in the probability simplex. We now show that student predictions trained via KD also exhibit structured bias:
@@
Let the teacher model $f$ exhibit structured prediction bias as defined in Def. 1, with $\{\mu_{k}\}_{k=1}^{K}$ and $\{r_{k}\}_{k=1}^{K}$. Assume the student model $f_{r}$ is trained via KD from $f$ using the loss $\lambda\mathcal{L}_{\text{CE}}+(1-\lambda)\mathcal{L}_{\text{KD}}$, and satisfies $\|f_{r}(x)-f^{*}(x)\|_{1}\leq\epsilon$ for all $x\in\mathcal{X}$, where $f^{*}(x)=\lambda y+(1-\lambda)f(x)$ and $y\in\{0,1\}^{C}$ denotes the label of $x$. Then student predictions $f_{r}(x)$ exhibit structured prediction bias as in Def. 1. Specifically, $\forall x\in\mathcal{X}$, there exists $k\in[K]$ such that:
@@
where the propagated centriod is defined as $\hat{\mu}_{k}(x)=\lambda y+(1-\lambda)\mu_{k}$, and the adjusted radius is $\hat{r}_{k}=(1-\lambda)r_{k}+\epsilon$. Since $y$ is one-hot and $\mu_{k}$ is fixed, the set of possible centers $\{\hat{\mu}_{k}(x)\}$ is finite and contained in $\Delta^{C-1}$, thus satisfying the condition of Def. 1 with at most $C\cdot K$ clusters.
@@
We defer the proof of Prop. 1 to Sec. A.2. Prop. 1 is also illustrated in Fig. 3-(left), where one of the teacher cluster, characterized by $(\mu_{k},r_{k})$, propagates to corresponding active student clusters with $(\hat{\mu}_{k},\hat{r}_{k})$ across three different class labels.
Empirical validation of Prop. 1.
@@
We train two models under a 1-shot setting (i.e., one labeled image per class): 1) a baseline model using $\mathcal{L}_{\text{CE}}$ on $\mathcal{D}^{(l)}$, and 2) a student model trained via KD with loss $\lambda\mathcal{L}_{\text{CE}}+(1-\lambda)\mathcal{L}_{\text{KD}}$ on $\mathcal{D}^{(l)}$ and $\mathcal{D}^{(u)}$. To assess bias propagation, we cluster teacher predictions on $\mathcal{D}^{(u)}$ via $k$-means [28], yielding centroids $\{\mu_{k}\}_{k=1}^{K}$. For each $x\in\mathcal{D}^{(u)}$, we identify its assigned teacher centroid $\mu_{k}$ and compute the corresponding propagated student centroid $\hat{\mu}_{k}=\lambda y+(1-\lambda)\mu_{k}$, where $y$ is the one-hot ground-truth label of $x$. We then measure the average $\ell_{2}$ distance between $f_{r}(x)$ and $\hat{\mu}_{k}$ for both models, using a threshold $\epsilon=0.3$. As shown in Fig. 4(c), the distilled student shows consistently lower distances than the baseline across varying $K$, validating Prop. 1. The number of active propagated centroids also closely matches that of the teacher, further supporting the structured prediction bias of student model.
Probabilistic Coreset (PCoreSet) selection.
@@
In this work, we interpret the structured prediction biases of teacher models as an inductive bias in the ActiveKD setting. During KD, the student model has access to labeled data and therefore does not merely mimic the teacher’s outputs, but learns from them under supervision. As a result, the teacher’s prediction structure acts as a prior that shapes the student’s hypothesis space. Our key insight is that when KD is effective (with bounded error $\epsilon$), the student inherits the teacher’s clustered structure in the probability space (Prop. 1). Samples that deviate from this structure are particularly informative, as they fall outside regions captured by the inductive bias. Rather than discarding this structure, we leverage it to guide active learning—selecting samples that challenge the current bias and expand the student’s predictive capacity.
@@
To this end, we propose Probabilistic Coreset (PCoreSet), a simple yet effective selection strategy inspired by coreset selection [76]. While conventional coreset methods aim to maximize coverage in the feature space, PCoreSet instead maximizes coverage in the probability simplex $\Delta^{C-1}$ by targeting underrepresented probability regions. This enables the student to inherit the teacher’s inductive bias more completely while actively exploring beyond it. Formally, given a labeled dataset $\mathcal{D}^{(l)}$, we greedily select $x^{*}\in\mathcal{D}^{(u)}$ as:
@@
where $d(x,x^{\prime}):=\|f_{r}(x)-f_{r}(x^{\prime})\|_{2}$ measures distance in the probability space. We then request the label $y^{*}$ for $x^{*}$ and update the labeled set: $\mathcal{D}^{(l)}\leftarrow\mathcal{D}^{(l)}\cup\{(x^{*},y^{*})\}$. PCoreSet thus generalizes coreset selection to the probability space, encouraging coverage of underrepresented predictions to better align with and extend the inductive bias learned via KD. See Alg. 2 for details.
@@ 4 Experiments
In this section, we evaluate the empirical effectiveness of our proposed methods. We present our experimental setups in Sec. 4.1 and the following empirical findings in Sec. 4.2:
  • [F1]: ActiveKD significantly enhances active learning performance by leveraging the knowledge of both zero- and few-shot VLM teachers.
  • [F2]: PCoreSet promotes diverse sampling in the probability space, allowing the student to both inherit and extend the inductive biases of teacher models, leading to improved performance under the proposed ActiveKD framework.
@@ 4.1 Experimental Setup
We evaluate our approach across 11 diverse datasets: generic image recognition benchmarks such as ImageNet [73] and Caltech101 [20], and fine-grained and domain-specific benchmarks such as StanfordCars [47], Flowers102 [63], FGVCAircraft [57], Food101 [9], SUN397 [91], DTD [15], EuroSAT [33], and UCF101 [80]. See an overview of datasets in Appendix C.
@@
We evaluate three distinct active learning (AL) frameworks:
  1. 1.
    No Distill: standard AL without knowledge distillation, where the model is trained solely on labeled data $\mathcal{D}^{(l)}$ using the cross-entropy loss $\mathcal{L}_{\text{CE}}$.
  2. 2.
    Zero-Shot: our proposed ActiveKD framework with a fixed zero-shot VLM teacher model [68] that provides soft targets for distillation on both $\mathcal{D}^{(l)}$ and $\mathcal{D}^{(u)}$.
  3. 3.
    Few-Shot: same as 2. Zero-Shot but with a few-shot VLM teacher model [77] that is fine-tuned on $\mathcal{D}^{(l)}$ with newly selected samples at each round.
@@
We compare our proposed method, PCoreSet, with five representative active selection strategies:
  1. 1.
    Random: randomly selects samples from the unlabeled pool $\mathcal{D}^{(u)}$.
  2. 2.
    Entropy [35]: selects samples with the highest prediction entropy.
  3. 3.
    Coreset [76]: maximizes sample diversity in the feature space.
  4. 4.
    BADGE [4]: A hybrid approach that combines uncertainty and diversity via gradient embeddings.
  5. 5.
    Class-Balanced [6]: prioritizes class diversity in the selected samples.
@@
All AL frameworks begin from a 1-shot setting, where one labeled example is provided per class across $C$ classes. We use a query size of $Q=C$ per round, with $R=16$ rounds for all datasets except ImageNet, which uses $R=8$. We adopt DHO [41] as the KD method in all experiments to leverage the unlabeled pool. For ImageNet, we use a self-supervised ResNet-50 [11] as the student model. For the other 10 datasets, we use ResNet-18 [29], MobileNetV2 [75], and TinyViT [89], all pretrained on ImageNet. We consistently use CLIP ResNet-50 [68] as the zero-shot teacher and CLAP [77] as the few-shot teacher. To reduce computational cost, we subsample the unlabeled ImageNet pool to 100,000 images. We exclude BADGE [4] on ImageNet due to memory limitations when computing gradients over its 1,000 classes. We report the mean performance with 95% confidence intervals across 5 random seeds. We defer further details to Appendix B, e.g., pseudocode for KD, training hyperparameters, and baseline implementations.
4.2 Experimental Results
@@
Table 1: Results on ImageNet and average across 10 datasets with different AL frameworks at the final AL round. We report average and 95% CI over 5 runs.
Strategy No Distill Zero-Shot Few-Shot
ImageNet Performance
Random 33.36 ± 0.45 60.69 ± 0.16 60.49 ± 0.12
Coreset [76] 26.61 ± 0.90 60.58 ± 0.07 59.01 ± 0.42
Entropy [35] 30.76 ± 0.24 60.43 ± 0.16 58.87 ± 0.14
ClassBalanced [6] 34.44 ± 0.60 61.07 ± 0.11 61.13 ± 0.07
PCoreSet (ours) 33.41 ± 0.41 61.16 ± 0.14 61.57 ± 0.24
Average over 10 Datasets
Random 63.10 ± 4.33 76.31 ± 0.88 77.48 ± 0.86
Coreset [76] 64.99 ± 3.01 76.20 ± 0.91 77.63 ± 0.57
Entropy [35] 64.06 ± 3.47 75.58 ± 0.86 77.47 ± 0.90
Badge [4] 63.14 ± 4.66 76.18 ± 0.74 77.50 ± 0.67
ClassBalanced [6] 63.57 ± 4.33 76.28 ± 0.87 77.56 ± 0.66
PCoreSet (ours) 64.94 ± 3.25 77.44 ± 0.84 78.81 ± 0.42
Figure 5: Results on ImageNet (Top) and average across 10 datasets (Bottom) over AL rounds. Zero- and Few-shot share the y-axis for comparison.
[F1-1]: Effectiveness of ActiveKD with zero-shot teachers.
We evaluate our ActiveKD framework using zero-shot VLM teachers. Table 1 shows that ActiveKD (Zero-Shot) consistently outperforms the baseline without distillation across all datasets and selection strategies. Zero-shot distillation improves accuracy by 27.33% on ImageNet with random selection (33.36% $\to$ 60.69%) and by 29.07% on average across strategies. This improvement extends across 10 datasets with an average gain of 12.21% (63.10% $\to$ 76.31%) for random selection, and 12.37% on average across all strategies, demonstrating the effectiveness of zero-shot distillation under limited supervision. Fig. 5 confirms these performance gains persist throughout the AL process.
Figure 6: Results of few-shot teachers on 11 datasets across rounds on ImageNet (Left) and averaged across 10 datasets (Right).
[F1-2]: Effectiveness of ActiveKD with few-shot teachers.
Few-shot distillation further enhances performance, as shown in Table 1, boosting accuracy by 0.41% on ImageNet and 1.37% across 10 datasets when paired with PCoreSet. While Entropy and Coreset show slight performance drops when transitioning from zero-shot to few-shot, indicating that teacher fine-tuning benefits vary by dataset and selection strategy, PCoreSet achieves the most substantial gains, demonstrating strong synergy with few-shot distillation. Fig. 5 confirms these improvements persist across most AL rounds, with exceptions only in early ImageNet rounds ($r<4$).
@@
Figure 7: Results on 10 datasets using ResNet-18 under zero-shot distillation across 16 rounds.
Figure 8: Results on 10 datasets using TinyViT under zero-shot distillation across 16 rounds.
To better assess the effectiveness of PCoreSet in ActiveKD with few-shot teachers, we track teacher performance across AL rounds. As shown in Fig. 6, PCoreSet consistently outperforms alternative selection strategies by enhancing few-shot teacher performance. This shows a key insight: structured prediction bias in the teacher propagates to the student during distillation. By selecting samples underrepresented in the student prediction space, PCoreSet also captures underrepresented modes in the teacher space. These selections not only boost student performance but also improve the teacher itself, forming a virtuous cycle of better selection and stronger distillation.
[F2-1]: Effectiveness of PCoreSet under ActiveKD.
We conduct an in-depth evaluation of PCoreSet under ActiveKD using ResNet-18 student and zero-shot teacher. As shown in Fig. 7, PCoreSet outperforms active selection baselines across all AL rounds and 10 datasets in the most settings. Notably, none of the widely used baselines, Entropy, Coreset, or BADGE, demonstrates consistent superiority over Random. In fact, Random surpasses these baselines on several datasets, including ImageNet (Table 1), FGVC Aricraft, and Food101, while remaining competitive on others. This unexpected effectiveness of Random underscores the complexity of developing effective active learning methods for KD. We further validate the generalizability of PCoreSet across different architecture families: TinyViT [89] (vision transformer) and MobileNetV2 [75] (efficient convolutional network). As shown in Fig. 8, PCoreSet outperforms baselines with TinyViT under zero-shot distillation in the most settings. MobileNetV2 exhibits similar trends (see Sec. D.1).
@@
Figure 9: KNN loss (L2) and cluster purity of PCoreSet and other selection strategies across 10 datasets.
[F2-2]: The acceleration of structured bias propagation of PCoreSet.

### v2

@@ 3.2 ActiveKD
In this paper, we propose the ActiveKD framework that leverages knowledge distillation into active learning. Our framework consists of two key components: 1) training the student model with zero-/few-shot teachers, and 2) performing sample selection using the distilled student model, potentially in collaboration with the teacher model. Specifically, we train the student model for each round using both supervised learning and knowledge distillation (Hinton et al., 2015; Kang et al., 2025):
@@
where the final loss is $\lambda\mathcal{L}_{\text{CE}}+(1-\lambda)\mathcal{L}_{\text{KD}}$, $\ell$ denotes $D_{\mathrm{KL}}$ represent cross-entropy and Kullback-Leibler divergence, respectively. For each round $r$, we use a mini-batch version of the above objective with stochastic gradient descent to optimize the parameters of $f_{r}$ by leveraging the teacher prediction $f(\cdot)$. See underlines in Algorithm˜1 for additional parts of ActiveKD.
@@
Figure 4: Visualization of prediction bias and its propagation. (a): hard prediction bias of the teacher; (b): soft prediction bias of the teacher (left) and the student (right) after distillation; and (c): the bias propagated from the teacher to the student, quantified by KNN loss ($\ell_{2}$) across different numbers of clusters ($K$).
VLMs such as CLIP ResNet-50 (Radford et al., 2021) inherently exhibit a class imbalance (Bang et al., 2024) due to their pre-training data distribution and language prompt design, as shown in Fig.˜4(a). Beyond class imbalance, we observe that teacher predictions form distinct clusters that occupy constrained regions of the probability simplex. As visualized in Fig.˜4(b)-(left), the violin plots show probability distributions of samples for given ground-truth (GT) classes, revealing how teacher predictions cluster in specific regions rather than uniformly covering the probability space, creating blind spots in the predictive capability of the model. This clustering behavior propagates to student models during knowledge distillation, as we observe that students inherit similar biased prediction patterns as shown in Fig.˜4(b)-(right). To formalize this, we first define the teacher’s structured prediction bias as follows:
@@
In other words, all teacher predictions lie within a finite union of $\ell_{2}$-balls in the probability simplex. We now show that student predictions trained via KD also exhibit similar structured bias:
@@
Let the teacher model $f$ exhibit structured prediction bias as defined in Definition˜1, with $\{\mu_{k}\}_{k=1}^{K}$ and $\{r_{k}\}_{k=1}^{K}$. Assume the student model $f_{r}$ is trained via KD from $f$ using the loss $\lambda\mathcal{L}_{\text{CE}}+(1-\lambda)\mathcal{L}_{\text{KD}}$, and satisfies $\|f_{r}(x)-f^{*}(x)\|_{1}\leq\epsilon$ for all $x\in\mathcal{X}$, where $f^{*}(x)=\lambda y+(1-\lambda)f(x)$ and $y\in\{0,1\}^{C}$ denotes the label of $x$. Then student predictions $f_{r}(x)$ also exhibit structured prediction bias defined in Definition˜1. Specifically, $\forall x\in\mathcal{X}$, there exists $k\in\{1,\ldots,K\}$ such that:
@@
where the propagated centriod is defined as $\hat{\mu}_{k}(x)=\lambda y+(1-\lambda)\mu_{k}$, and the adjusted radius is $\hat{r}_{k}=(1-\lambda)r_{k}+\epsilon$. Since $y$ is one-hot and $\mu_{k}$ is fixed, the set of possible centers $\{\hat{\mu}_{k}(x)\}$ is finite and contained in $\Delta^{C-1}$, thus satisfying the condition of Definition˜1 with at most $C\cdot K$ clusters.
@@
We defer the proof of Proposition˜1 to Section˜A.2. Proposition˜1 is also illustrated in Fig.˜3-(left), where one of the teacher cluster, characterized by $(\mu_{k},r_{k})$, propagates to corresponding active student clusters with $(\hat{\mu}_{k},\hat{r}_{k})$ across three different class labels.
Empirical validation of Proposition˜1.
@@
We train two models under a 1-shot setting (i.e., one labeled image per class): 1) a baseline model using $\mathcal{L}_{\text{CE}}$ on $\mathcal{D}^{(l)}$, and 2) a student model trained via KD with loss $\lambda\mathcal{L}_{\text{CE}}+(1-\lambda)\mathcal{L}_{\text{KD}}$ on $\mathcal{D}^{(l)}$ and $\mathcal{D}^{(u)}$. To assess bias propagation, we cluster teacher predictions on $\mathcal{D}^{(u)}$ via $k$-means (Hartigan & Wong, 1979), yielding centroids $\{\mu_{k}\}_{k=1}^{K}$. For each $x\in\mathcal{D}^{(u)}$, we identify its assigned teacher centroid $\mu_{k}$ and compute the corresponding propagated student centroid $\hat{\mu}_{k}=\lambda y+(1-\lambda)\mu_{k}$, where $y$ is the one-hot GT label of $x$. We then measure the average $\ell_{2}$ distance between $f_{r}(x)$ and $\hat{\mu}_{k}$ for both models, using a threshold $\epsilon=0.3$. As shown in Fig.˜4(c), the distilled student shows consistently lower distances than the baseline across varying $K$, validating Proposition˜1. The number of active propagated centroids also closely matches that of the teacher, further supporting the existence of structured prediction bias of student model.
3.3 Probabilistic Coreset (PCoreSet)
Motivation.
@@ 3.3 Probabilistic Coreset (PCoreSet)
During KD, the student model leverages both labeled data and teacher guidance, learning to generalize rather than merely mimicking teacher outputs. As a result, the prediction structure of the teacher acts as a prior that shapes the hypothesis space of the student. Our key insight is that when KD is effective (with bounded error $\epsilon$), the student inherits the clustered structure of teacher in the probability space (Proposition˜1). Samples that deviate from this structure are particularly informative, as they fall outside regions captured by the inductive bias. Rather than discarding this structure, we leverage it to guide active learning: selecting samples that deviate from the established structure and expand the student’s predictive capacity.
PCoreSet selection.
@@
To this end, we propose Probabilistic Coreset (PCoreSet), a simple yet effective selection strategy inspired by coreset selection (Sener & Savarese, 2017). While conventional coreset methods aim to maximize coverage in the feature space, PCoreSet instead maximizes coverage in the probability simplex $\Delta^{C-1}$ by targeting underrepresented probability regions. This enables the student to inherit the teacher’s inductive bias more completely while actively exploring beyond it. Formally, given a labeled dataset $\mathcal{D}^{(l)}$, we greedily select $x^{*}\in\mathcal{D}^{(u)}$ as:
@@
where $d(x,x^{\prime}):=\|f_{r}(x)-f_{r}(x^{\prime})\|_{2}$ measures distance in the probability space. We then request the label $y^{*}$ for $x^{*}$ and update the labeled set: $\mathcal{D}^{(l)}\leftarrow\mathcal{D}^{(l)}\cup\{(x^{*},y^{*})\}$. See Algorithm˜2 for details.
Computational complexity.
Note that the computational complexity of PCoreSet is $\mathcal{O}(C\cdot M\cdot N)$, whereas the feature-space coreset (Sener & Savarese, 2017) has complexity $\mathcal{O}(H\cdot M\cdot N)$, where $H$ is the feature dimensionality. Thus, in most cases ($C\ll H$), PCoreSet is more efficient.
@@ 4.1 Experimental Setup
We evaluate our approach across 11 diverse datasets: generic image recognition benchmarks (Russakovsky et al., 2015; Fei-Fei et al., 2004), and fine-grained and domain-specific benchmarks (Krause et al., 2013; Nilsback & Zisserman, 2008; Maji et al., 2013; Bossard et al., 2014; Xiao et al., 2010; Cimpoi et al., 2014; Helber et al., 2019; Soomro et al., 2012). To reduce computational cost, we subsample the unlabeled ImageNet (Russakovsky et al., 2015) pool to 100,000 images. See an overview of datasets in Appendix˜C.
@@
1) No Distill: standard AL without knowledge distillation, where the model is trained only on labeled data $\mathcal{D}^{(l)}$ using cross-entropy loss $\mathcal{L}_{\text{CE}}$; 2) ActiveKD (Zero-Shot): our proposed ActiveKD framework with a fixed zero-shot VLM teacher (Radford et al., 2021) that provides soft targets for distillation on both $\mathcal{D}^{(l)}$ and $\mathcal{D}^{(u)}$; and 3) ActiveKD (Few-Shot): as in 2) ActiveKD (Zero-Shot) but with a few-shot VLM teacher (Silva-Rodriguez et al., 2024) fine-tuned on $\mathcal{D}^{(l)}$ using both the newly selected samples and existing labeled data.
@@
We compare PCoreSet with five baselines. 1) Random: uniform sampling from the unlabeled pool $\mathcal{D}^{(u)}$; 2) Entropy (Holub et al., 2008): selects points with highest predictive entropy; 3) CoreSet (Sener & Savarese, 2017): maximizes diversity in feature space; 4) BADGE (Ash et al., 2019): combines uncertainty and diversity via gradient embeddings; and 5) Class-Balanced (Bang et al., 2024): promotes class diversity in the queried set.
@@
All AL methods begin from a 1-shot setting, where one labeled example is provided per class across $C$ classes. We use a query size of $Q=C$ per round, with $R=16$ rounds for all datasets except ImageNet, which uses $R=8$. We adopt DHO (Kang et al., 2025) as the KD method in all experiments to leverage the unlabeled pool. For ImageNet, we use a self-supervised ResNet-50 (Caron et al., 2021) as the student model. For the other 10 datasets, we use ResNet-18 (He et al., 2016), MobileNetV2 (Sandler et al., 2018), TinyViT (Wu et al., 2022), and ViT-B/16 (Dosovitskiy et al., 2020), all pretrained on ImageNet. We use CLIP ResNet-50 (Radford et al., 2021), OpenCLIP (Cherti et al., 2023), and ViT-L/14 as the zero-shot teacher and CLAP (Silva-Rodriguez et al., 2024) as the few-shot teacher. We exclude BADGE (Ash et al., 2019) on ImageNet due to memory limitations when computing gradients over its 1,000 classes. We report the mean performance with 95% confidence intervals across 5 random seeds. We defer further details to Appendix˜B.
4.2 Main Results
@@
Table 1: Results on ImageNet and the average over 10 datasets under different AL frameworks at the final AL round. We report the mean and 95% CI over five runs; values in parentheses denote the gain over No Distill. ActiveKD consistently improves over No Distill across all strategies and datasets.
ImageNet (ResNet-50 student & ResNet-50 teacher) Avg. over 10 Datasets (ResNet-18 student & ResNet-50 teacher)
Methods No Distill ActiveKD (Zero-Shot) ActiveKD (Few-Shot) No Distill ActiveKD (Zero-Shot) ActiveKD (Few-Shot)
[…]
Avg. over Methods 31.72 60.79 (+29.07) 60.21 (+28.50) 63.97 76.33 (+12.37) 77.74 (+13.78)
Figure 5: Results on ImageNet and the average over 8 or 10 datasets across 5 student and 3 teacher architectures under ActiveKD (Zero-Shot) with either 8 or 16 rounds. We report the mean and 95% CI over five runs. PCoreSet achieves the best performance in 64/73 settings ($\approx$87.7%)
@@
Figure 6: Results on 10 datasets using ResNet-18 with ResNet-50 teacher under ActiveKD (Zero-Shot).
Effectiveness of ActiveKD (Zero-Shot).
We first evaluate ActiveKD with ResNet-18/ResNet-50 students and a ResNet-50 zero-shot teacher. Tab.˜1 shows that ActiveKD (Zero-Shot) consistently improves No Distill across all datasets and selection strategies: e.g., it improves accuracy by +27.33% on ImageNet with Random selection (33.36% $\to$ 60.69%) and by +29.07% on average across strategies. This improvement extends across 10 datasets with an average gain of +13.21% (63.10% $\to$ 76.31%) for Random selection, and +12.37% on average across all strategies. These substantial gains demonstrates the effectiveness of our proposed ActiveKD under limited supervision.
Effectiveness of ActiveKD (Few-Shot).
ActiveKD (Few-Shot) also consistently improves over No Distill across all datasets and strategies; relative to ActiveKD (Zero-Shot) it yields +1.17% for Random and +1.41% on average across the 10 datasets, demonstrating the benefit of strong few-shot teachers. However, the improvements of ActiveKD (Few-Shot) are not consistent on ImageNet; e.g., the gains are only +0.06% with ClassBalanced and +0.41% with PCoreSet. We attribute this to the large number of classes in ImageNet (C=1000C{=}1000), which increases labeled samples per round, reducing the added value of teacher signals compared with other 10 datasets having $C\approx 100$. For the same reason, Random often outperforms Entropy/Coreset on ImageNet—also noted by Emam et al. (2021); Bang et al. (2024)—making Random a strong baseline.
Effectiveness of PCoreSet under ActiveKD (Zero-Shot).
We next evaluate PCoreSet under ActiveKD (Zero-Shot). Fig.˜5 reports results for different selection strategies on ImageNet and on the averages over 10 or 8 datasets across 8 or 16 AL rounds. We evaluate across a comprehensive set of 73 experimental settings, calculated as $({\color[rgb]{0,0,0.8}8}-1)\times{\color[rgb]{0.8,0,0}4}+({\color[rgb]{0,0,0.8}16}-1)\times{\color[rgb]{0.8,0,0}3}$. This computation accounts for 8-round and 16-round active learning experiments, where we exclude the first round since all models begin with identical training sets, making strategy comparisons meaningless at that round. The factors 4 and 3 represent the number of experimental configurations for the 8-round and 16-round setups respectively, encompassing both ImageNet-specific results and cross-dataset averages under various conditions. Across these 73 settings, PCoreSet ranks first in 64 (64/73 $\approx$ 87.7%), always achieving the best performance except for first 2 AL rounds, showing that it becomes increasingly effective as active learning progresses.
In-depth comparison.
Fig.˜6 presents an in-depth evaluation of selection strategies under ActiveKD (Zero-Shot) across 10 datasets, using a ResNet-18 student and a ResNet-50 teacher. We observe that PCoreSet achieves the best performance or is at least on par with the best. Similar trends appear in other in-depth evaluations across different settings; see Appendix˜D for details.
@@
Figure 7: KNN loss (L2) and cluster purity of PCoreSet and other selection strategies across 10 datasets, and results of few-shot teachers on 11 datasets across rounds.
4.3 Analysis
The acceleration of structured bias propagation.